# Mapping a Movement: Exploring a Proposed Police Training Facility in Atlanta and the Stop Cop City Movement through Online Maps

### v1

@@ Abstract.
ccs: Human-centered computing Empirical studies in HCIccs: Human-centered computing Geographic visualization
Figure 1. A sample of maps retrieved from online sources that describe aspects of a proposed police training facility (colloquially known as Cop City) in Atlanta, Georgia, from proponents of the facility, the mass media, and activists who oppose the plan.
[…]
Mosaic of maps related to the Proposed Police Training Facility
@@ 1.1. Background and Motivation
The cartographic process and map products are affected by the mapmaker’s purpose and can be used to execute a strategy (Eichberger, 2019; Mackaness, 1995). For example, maps that explain human-environmental issues such as fracking may be viewed differently depending on the stakeholders presenting the information and the information being included or omitted. For this reason, it is valuable to study how maps are made, what they contain, and what strategies and technology are used, especially when describing a site undergoing a contentious project. This form of analysis can capture the variation in technologies and strategies used by various stakeholders who support the project, oppose the project, or appear to remain neutral, and how these groups use technology to convey their message.
@@ 2.1. The Proposed Police Training Facility
The proposal was developed amidst mass anti-police and abolitionist uprisings in Atlanta and across the country in 2020. This was driven nationally by the Minneapolis Police murder of George Floyd in May 2020 and driven further locally by the Atlanta Police murder of Rayshard Brooks in June 2020.
@@
The site of the proposed facility has raised concerns for several reasons. The site is located in the South River Forest, which is part of Atlanta’s four key sources of tree canopy, known locally as the “lungs of Atlanta”. These forested areas are crucial for mitigating harmful impacts of climate change in Atlanta such as extreme heat, flooding, and air pollution (Muse et al., 2022; Pallathadka et al., 2022). Atlanta’s ample tree canopy is also endangered (Evans et al., 2023) and increasing development in the area has resulted in a reduction in the canopy. Furthermore, the site is known for its history as a prison farm, where incarcerated individuals, predominately Black men, were forced to work from the 1920s to as recently as the 1990s (Newman et al., 2022; Mowatt, 2023; Gant, 2022). Prior to the construction of the prison farm, the Muscogee Creek peoples were expelled from the land during the Trail of Tears in the late 1800s. The project raises concerns for continuing this legacy of exploitation of marginalized peoples, particularly of Black and indigenous populations which have been historically exploited, expelled from the area, and are at a higher risk of violence and murder by police (Allard et al., 2010; Byrd, 2017). These concerns are exacerbated by timing of the proposal for the site. Because the proposal came amidst the anti-police and abolitionist uprisings, the SCC movement purports that the project will likely lead to increased police suppression of social movements [^1] Today News: https://grassrootsthinking.com/2023/01/26/from-atlanta-to-memphis-we-demand-justice-stop-cop-city/, and an increase in militaristic and deadly practices in everyday policing of marginalized Black communities.
@@
The movement has garnered national attention due to key escalations that took place in 2023; In January 2023, Georgia Bureau of Investigation officers killed activist Manuel Paez Teran (also known as Tortuguita) in the South River Forest (Mowatt, 2023). Subsequently, activists responded with street protests, vigils, and concerts. These events, collectively, resulted in the arrests of 61 individuals, who were subsequently charged with RICO [^2] RICO stands for the Racketeer Influenced and Corrupt Organizations Act. violations and domestic terrorism charges. In the summer of 2023, activists launched the “Let Atlanta Decide” effort, a campaign to force a legal referendum to allow Atlantans to vote to repeal the 2021 ordinance that authorized the city to lease the project site’s land. “Let Atlanta Decide” organizers claim that the effort garnered over 100,000 signatures from Atlanta City Residents.
@@
Through our review of academic literature (Akbar, 2023), news media coverage (Brown, 2023; Buchheit, 2023; Pratt, 2023), statements from grassroots organizations  [^3] Community Movement Builders, Stop Cop City: http://communitymovementbuilders.org/stop-cop-city/  [^4] Stop Cop City Solidarity: http://stopcopcitysolidarity.org/about  [^5] Gen-Z for Change, Stop Cop City: http://stopcopcity.net/ in opposition to the facility, and video of public comment at Atlanta City Council Meetings (Council, [n. d.]; jun, 2023), we identified key concerns about the facility. These include issues on the topics of environmental justice and environmental preservation, racial and economic justice, rights and well-being of local residents, police violence and over-policing, indigenous land protection, expenditure of funds collected from taxes, legal issues around land ownership, and voting rights. We organize the narrative of our results around these key concerns and how they are reflected in maps.
@@ 2.2. Critical Cartography
Maps are not objective representations. Instead, they are highly stylized by the mapmaker, technologies, and systems the mapmaker uses. Critical cartography questions the information, visual communication methods, reasoning and strategies behind the display, and the message that map readers receive (Harley, 1989). A critical analysis of maps can highlight the needs of those who have little power in map construction (Crampton, 2001). For example, the question of who and what is being excluded from a map is important because maps can ‘mask or hide’ an interest (Wood and Fels, 1992; Eichberger, 2019). Further, as Fish and Kreitzberg (2023) find in an analysis of map use (or lack thereof) in far right news media publications, the choice to use maps or not is a subjective decision that can be influenced by political motivations. Furthermore, many works within GIS have shown that usability challenges with mapmaking software and accessibility to such software impacts map production, especially for members of the public (Dransch, 2001; Rzeszewski and Kotus, 2019; Czepkiewicz et al., 2016; Jankowski et al., 2019). When given access to mapmaking technologies, individuals have expanded capabilities to create maps that fit their needs, although this capability is limited to what is permitted within the software (Dransch, 2001), as well as hardware constraints, time constraints, and access to high quality data.
@@
Public participation GIS is the use of GIS to broaden public involvement in policymaking and/or leveraging GIS to promote the goals of NGOs, grassroots groups, and community-based organizations. Prior literature on public participation GIS provides further insight on critical cartography in relation to power and silencing. For instance, Cidell (2008), conducts a participatory GIS study of residents’ challenges to sound maps through analysis of public comment recordings. They find that public comment enabled residents to challenge how data on ambient sound was collected and portrayed (as contour lines) and suggested enhanced resident involvement in the cartographic process. One critique of public participation GIS levied by Sieber is the following: “GIS distracts grassroots groups and others from proven activist strategies such as protest and retreats from questioning the general framework of policymaking and distribution of power”(Sieber, 2006, p. 491). Sieber argues that despite the growing ubiquity of GIS systems, the use of GIS technologies in grassroots organizations can lead to a false illusion of control, rather than using proven strategies to disrupt the actual dynamics of control and power. Despite this, Sieber points out that analysis of conventional models of participatory GIS do not include oppositional forms, such as uses of GIS to plan protests or riots, or other uses of GIS that explicitly challenge authority. Our work builds on this criticism of public participation GIS and analyzes oppositional use of GIS, while contrasting to institutionally aligned uses of GIS technology.
@@ 3.1.2. Exclusion Criteria
Before thematic coding all maps collected were manually screened for relevance and geographical accuracy to the APF site and/or Stop Cop City movement by both authors. After the initial data collection and screening, we applied additional exclusion criteria. We excluded maps that were used artistically as part of a flyer or a graphic. Unless the underlying map or the annotations themselves included unique information, we did not include these graphics. The same map image was frequently reposted by multiple sources, and we removed duplicated maps from our map corpus. Maps of the site that were created before the facility was proposed were rare, and were also not used in the analysis.
The data acquisition process resulted in a total of 45 unique map images, including annotated aerial photos (i.e., orthorectified imagery with graphics or freehand mouse-drawn drawings), site plans furnished by local government and contracted planners, architects, and civil engineers. We retrieved three maps from the EIS documents. We reduced this set of maps to 32 unique maps after applying the exclusion criteria described above. The full corpus of 45 maps (including a map found outside the search) and a brief description of the maps found in the EIS documents is included in the supplementary material. Table 1 denotes the full set of 32 maps used for keyword analysis. Most of the 32 maps we retrieved, including two interactive maps (rows 21 and 32 in Table 1), were sourced from news articles.
@@ 4. Results
No. Publication & Author (Hyperlinked) Data Sources, Software, Primary Sources Pub. Date
1 Terracon (page 91) Various sources such as USGS topographic maps and aerial imagery 4-22-21
2 Defend the Atlanta Forest (Twitter @$defendatlforest$) None stated 5-13-21
3,4 Mainlinezine, By Wayne Butler USGS StreamStats, APF Data 7-14-21
5 Urbanize Atlanta, By Josh Green Atlanta Police Foundation 8-25-21
6 Atlanta Journal Constitution, By Anjali Huynh Atlanta Police Foundation 8-10-21
7 The Appeal, By Aja Arnold None stated 12-8-21
8, 9 Terracon, Atlanta Community Press Collective, by Lily Ponitz. (Reproduction of maps by) Terracon 1-23-22
10-12 Atlanta Community Press Collective (Twitter @$atlanta$_$press$) Dekalb County Property Information Website, Google Sheets 5-11-22
13 Capital B News, By Adam Mahoney and Adjoa Danso OpenStreetMap, Datawrapper 6-6-22
14 Defend the Atlanta Forest Google Maps 7-16-22
15 The Atlanta Journal-Constitution, By Tyler Estep, Map: Hannah Ziegler OpenStreetMap, Datawrapper 8-12-22
16, 17 CNN, By Christina Maxouris, Map: Renee Rigdon South River Forest Coalition, Atlanta Police Foundation, Google 9-24-22
18 Axios, By Emma Hurt City of Atlanta 2-1-23
19 Fleximillien, (Twitter @$thatabananaor$) WABE.org, Atlanta Regional Commission, U.S. Census Bureau 2-6-23
20 Saporta Report, By John Ruch Affidavit of R. Baskin, Atlanta Police Foundation 3-1-23
21 Atlanta Community Press Collective OpenStreetMap, Datawrapper 3-4-23
22 National Public Radio, By Bill Chappell City of Atlanta 3-7-23
23 Inside Climate News, By Victoria St. Martin, Map: Paul Horn Atl. Community Press Collective, Atl. Police Foundation, ESRI 3-8-23
24 Reddit.com/r/Weelaunee, (u/n0noTAGAinnxw4Yn3wp7) CAD, none stated 4-24-23
25 Saporta Report by John Rush South River Watershed Alliance 5-1-23
26 Reddit.com/r/Weelaunee, (u/n0noTAGAinnxw4Yn3wp7) Georgia Public Safety Training Center 5-20-23
27 Mapping Atlanta, By Taylor Shelton QGIS 5-30-23
28 The Xylom, By Alex Ip (@$AlexIp718$) OpenStreetMap, Datawrapper 8-2-23
29, 30 Anti-Racist South, (Twitter @$antiracistsouth$) Environmental Defense Fund, Texas A&M Univ., Darkhorse Analytics, MapBox, OpenStreetMap 10-2-23
@@ 28 The Xylom, By Alex Ip (@$AlexIp718$) OpenStreetMap, Datawrapper 8-2-23
32 StopCopCitySolidarity.org via RAM INC (Twitter @$resistabolish$) Leaflet, uMap, Django, OpenStreetMap N.D.
@@ 4.2.2. Elements Related to Socioeconomic Issues
Figure 7. Maps depicting analysis of socioeconomic data of the neighborhoods relevant to the site. Map A includes the highest (orange) and lowest (blue) income areas in the Atlanta metro area. using census tracts Images B and C depict levels of climate risk with very high risk in dark blue and low risk in cream. Map B depicts the county where many supporters of the facility live, while map C depicts the county where the facility is slated to be built. Maps A-C correspond to maps 19, 30, and 29, respectively, in Table 1.
@@ 4.2.4. Elements Related to Incarceration, Policing, Violence, and Danger
Narratives about abolition and over-policing in the SCC movement relate to physical features on, near, or in site plans, that signify incarceration and policing. Many maps point out landmarks related to incarceration (n=11). The Old Atlanta Prison Farm was the most prominently referenced landmark, and appeared on six maps, including maps by Terracon, the mass media, and the opposition. Maps also referenced the Youth Detention Center (n=3), which is also labeled as Prison for Children or Youth Prison (n=1 each) (Figures 9A and 9B). These two maps appear to have used other maps (such as Google Maps in Figure 9A; origin unknown in Figure 9B) for their arguments. The mapmapers appear to have annotated these images using other software that has free drawing features instead of making map labels, polygons, and annotations within the mapmaking software.

### v2

@@ Abstract.
$c$s: Human-centered computing Empirical studies in HCIccs: Human-centered computing Geographic visualization
Figure 1. A sample of maps retrieved from online sources that describe aspects of a proposed police training facility (colloquially known as Cop City) in Atlanta, Georgia, from proponents of the facility, the mass media, and activists who oppose the plan.:
@@ 1.1. Background and Motivation
It is well-known that the cartographic process and map products are affected by the mapmaker’s purpose and can be used to execute a strategy (Eichberger, 2019; Mackaness, 1995). For example, maps that explain human-environmental issues such as fracking may be viewed differently depending on the stakeholders presenting the information and the information being included or omitted. For this reason, it is valuable to study how maps are made, what they contain, and what strategies and technology are used, especially when describing a site undergoing a contentious project. This form of analysis can capture the variation in technologies and strategies used by various stakeholders who support the project, oppose the project, or appear to remain neutral, and how these groups use technology to convey their message.
@@ 2.1. The Proposed Police Training Facility
The proposal was developed amid mass anti-police and abolitionist uprisings in Atlanta and across the country in 2020. This was driven nationally by the Minneapolis Police murder of George Floyd in May 2020 and driven further locally by the Atlanta Police murder of Rayshard Brooks in June 2020.
@@
The site of the proposed facility has raised concerns for several reasons. The site is located in the South River Forest, which is part of Atlanta’s four key sources of tree canopy, known locally as the “lungs of Atlanta”. These forested areas are crucial for mitigating harmful impacts of climate change in Atlanta such as extreme heat, flooding, and air pollution (Muse et al., 2022; Pallathadka et al., 2022). Atlanta’s ample tree canopy is also endangered (Evans et al., 2023) and increasing development in the area has resulted in a reduction in the canopy. Furthermore, the site is known for its history as a prison farm, where incarcerated individuals, predominately Black men, were forced to work from the 1920s to as recently as the 1990s (Newman et al., 2022; Mowatt, 2023; Gant, 2022). Prior to the construction of the prison farm, the Muscogee Creek peoples were expelled from the land during the Trail of Tears in the late 1800s. The project raises concerns for continuing this legacy of exploitation of marginalized peoples, particularly of Black and indigenous populations which have been historically exploited, expelled from the area, and are at a higher risk of violence and murder by police (Allard et al., 2010; Byrd, 2017). These concerns are exacerbated by timing of the proposal for the site. Because the proposal came amid the anti-police and abolitionist uprisings, the SCC movement purports that the project will likely lead to increased police suppression of social movements, and an increase in militaristic and deadly practices in everyday policing of marginalized Black communities.
@@
The movement has garnered national attention due to key escalations that took place in 2023; In January 2023, Georgia Bureau of Investigation officers killed activist Manuel Paez Teran (also known as Tortuguita) in the South River Forest (Mowatt, 2023). Subsequently, activists responded with street protests, vigils, and concerts. These events, collectively, resulted in the arrests of 61 individuals, who were subsequently charged with RICO [^1] RICO stands for the Racketeer Influenced and Corrupt Organizations Act. violations and domestic terrorism charges. In summer 2023, activists launched the “Let Atlanta Decide” effort, a campaign to force a legal referendum to allow Atlantans to vote to repeal the 2021 ordinance that authorized the city to lease the project site’s land. “Let Atlanta Decide” organizers claim that the effort garnered over 100,000 signatures from Atlanta City Residents.
@@
Through our review of academic literature (Akbar, 2023), news media coverage (Brown, 2023; Buchheit, 2023; Pratt, 2023), statements from grassroots organizations in opposition to the facility, and video of public comment at Atlanta City Council Meetings (Council, [n. d.]; jun, 2023), we identified key concerns about the facility. These include issues on the topics of environmental justice and environmental preservation, racial and economic justice, rights and well-being of local residents, police violence and over-policing, indigenous land protection, expenditure of funds collected from taxes, legal issues around land ownership, and voting rights. We organize the narrative of our results around these key concerns and how they are reflected in maps.
@@ 2.2. Critical Cartography
Maps are not objective representations. Instead, they are highly stylized by the mapmaker, technologies, and systems the mapmaker uses. Critical cartography questions the information, visual communication methods, reasoning and strategies behind the display, and the message that map readers receive (Harley, 1989). A critical analysis of maps can highlight the needs of those who have little power in map construction (Crampton, 2001). For example, the question of who and what is being excluded from a map is important because maps can ‘mask or hide’ an interest (Wood and Fels, 1992; Eichberger, 2019). Further, as Fish and Kreitzberg (2023) find in an analysis of map use (or lack thereof) in far right news media publications, the choice to use maps or not is a subjective decision that can be influenced by political motivations. Furthermore, many works within GIS have shown that usability challenges with mapmaking software and accessibility; in turn, software impacts map production, especially for members of the public (Dransch, 2001; Rzeszewski and Kotus, 2019; Czepkiewicz et al., 2016; Jankowski et al., 2019). When given access to mapmaking technologies, individuals have expanded capabilities to create maps that fit their needs, although this capability is limited to what is permitted within the software (Dransch, 2001), as well as hardware constraints, time constraints, and access to high quality data.
@@
Public participation GIS can be describe as the use of GIS to broaden public involvement in policymaking and/or leveraging GIS to promote the goals of NGOs, grassroots groups, and community-based organizations. Prior literature on public participation GIS provides further insight on critical cartography in relation to power and silencing. For instance, Cidell (2008), conducts a participatory GIS study of residents’ challenges to sound maps through analysis of public comment recordings. They find that public comment enabled residents to challenge how data on ambient sound was collected and portrayed (as contour lines) and suggested enhanced resident involvement in the cartographic process. One critique of public participation GIS levied by Sieber is the following: “GIS distracts grassroots groups and others from proven activist strategies such as protest and retreats from questioning the general framework of policymaking and distribution of power”(Sieber, 2006, p. 491). Sieber argues that despite the growing ubiquity of GIS systems, the use of GIS technologies in grassroots organizations can lead to a false illusion of control, rather than using proven strategies to disrupt the actual dynamics of control and power. Despite this, Sieber points out that analysis of conventional models of participatory GIS do not include oppositional forms, such as uses of GIS to plan protests or riots, or other uses of GIS that explicitly challenge authority. Our work builds on this criticism of public participation GIS and analyzes oppositional use of GIS, while contrasting to institutionally aligned uses of GIS technology.
@@ 3.1.2. Exclusion Criteria
Before thematic coding all maps collected were screened by both authors to ensure all data included information about the proposed APF facility site and its geography. Following initial data collection, we excluded maps that were used artistically as part of a flyer or a graphic. Unless the underlying map or the annotations themselves included unique information, we did not include these graphics. The same map image was frequently reposted by multiple sources, and we removed duplicated maps from our map corpus. Maps of the site that were created before the facility was proposed were rare, and were also not used in the analysis.
The data acquisition process resulted in a total of 45 unique map images, including annotated aerial photos (i.e., orthorectified imagery with graphics or freehand mouse-drawn drawings), site plans furnished by local government and contracted planners, architects, and civil engineers. We retrieved three maps from the EIS documents. We reduced this set of maps to 32 unique maps after applying the exclusion criteria described above. The full corpus of 45 maps (including a map found outside the search) and a brief description of the maps found in the EIS documents is included in the supplementary materials. Table 1 denotes the full set of 32 maps used for keyword analysis. Most of the 32 maps we retrieved, including two interactive maps (rows 21 and 32 in Table 1), were sourced from news articles.
@@ 4. Results
No. Publication & Author Sources and Software Date
1 Terracon (page 91) Various sources such as USGS topographic maps and aerial imagery Apr 22, 2021
2 Defend the Atlanta Forest (Twitter @defendatlforest@defendatlforest) None stated May 13, 2021
3,4 Mainlinezine, By Wayne Butler USGS StreamStats, APF Data Jul 14, 2021
5 Urbanize Atlanta, By Josh Green Atlanta Police Foundation Aug 25, 2021
6 Atlanta Journal Constitution, By Anjali Huynh Atlanta Police Foundation Aug 10, 2021
7 The Appeal, By Aja Arnold None stated Dec 08, 2021
8, 9 Terracon, Atlanta Community Press Collective, by Lily Ponitz (Reproduction of maps by) Terracon Jan 23, 2022
10-12 Atlanta Community Press Collective (Twitter @atlanta@atlanta_presspress) Dekalb County Property Information Website, Google Sheets May 11, 2022
13 Capital B News, By Adam Mahoney and Adjoa Danso OpenStreetMap, Datawrapper Jun 6, 2022
14 Defend the Atlanta Forest Google Maps Jul 16, 2022
15 The Atlanta Journal-Constitution, By Tyler Estep, Map: Hannah Ziegler OpenStreetMap, Datawrapper Aug 12, 2022
16, 17 CNN, By Christina Maxouris, Map: Renee Rigdon South River Forest Coalition, Atlanta Police Foundation, Google Sep 24, 2022
18 Axios, By Emma Hurt City of Atlanta Feb 1, 2023
19 Fleximillien, (Twitter @thatabananaor@thatabananaor) WABE.org, Atlanta Regional Commission, U.S. Census Bureau Feb 6, 2023
20 Saporta Report, By John Ruch Affidavit of R. Baskin, Atlanta Police Foundation Mar 1, 2023
21 Atlanta Community Press Collective OpenStreetMap, Datawrapper Mar 4, 2023
22 National Public Radio, By Bill Chappell City of Atlanta Mar 7, 2023
23 Inside Climate News, By Victoria St. Martin, Map: Paul Horn Atlanta Community Press Collective, Atlanta Police Foundation, ESRI Mar 8, 2023
24 Reddit.com/r/Weelaunee, (u/n0noTAGAinnxw4Yn3wp7) CAD, none stated Apr 24, 2023
25 Saporta Report by John Rush South River Watershed Alliance May 1, 2023
26 Reddit.com/r/Weelaunee, (u/n0noTAGAinnxw4Yn3wp7) Georgia Public Safety Training Center May 20 2023
27 Mapping Atlanta, By Taylor Shelton QGIS May 30, 2023
28 The Xylom, By Alex Ip (@AlexIp718@AlexIp718) OpenStreetMap, Datawrapper Aug 2, 2023
29, 30 Anti-Racist South, (Twitter @antiracistsouth@antiracistsouth) Environmental Defense Fund, Texas A&M University, Darkhorse Analytics, MapBox, OpenStreetMap Oct 2, 2023
@@ 28 The Xylom, By Alex Ip (@AlexIp718@AlexIp718) OpenStreetMap, Datawrapper Aug 2, 2023
32 StopCopCitySolidarity.org via RAM INC (Twitter @resistabolish@resistabolish) Leaflet, uMap, Django, OpenStreetMap N.D.
@@ 4.2.2. Elements Related to Socioeconomic Issues
Figure 7. Maps depicting analysis of socioeconomic data of the neighborhoods relevant to the site. Map A includes the highest (orange) and lowest (blue) income areas in the Atlanta metro area using census tracts. Images B and C depict levels of climate risk with very high risk in dark blue and low risk in cream. Map B depicts the county where many supporters of the facility live, while map C depicts the county where the facility is slated to be built. Maps A-C correspond to maps 19, 30, and 29, respectively, in Table 1.
@@ 4.2.4. Elements Related to Incarceration, Policing, Violence, and Danger
Narratives about abolition and over-policing in the SCC movement relate to physical features on, near, or in site plans, that signify incarceration and policing. Many maps point out landmarks related to incarceration (n=11). The Old Atlanta Prison Farm was the most prominently referenced landmark, and appeared on six maps, including maps by Terracon, the mass media, and the opposition. Maps also referenced the Youth Detention Center (n=3), which is also labeled as Prison for Children or Youth Prison (n=1 each) (Figures 9A and 9B). These two maps appear to have used other maps (such as Google Maps in Figure 9A; origin unknown in Figure 9B) for their arguments. The mapmakers appear to have annotated these images using other software that has free drawing features instead of making map labels, polygons, and annotations within the mapmaking software.